# Iterative Reasoning Preference Optimization

### v2

@@ 3.2 ARC-Challenge Task
To test reasoning capabilities outside of mathematics, we employ ARC (Clark et al., 2018) which covers multiple science subjects. The training dataset contains 7.7k multiple-choice questions split into easy and challenge sets. We report results on the ARC-Challenge test set which has 1172 examples. There is no gold chain-of-thought reasoning provided for training examples in this task, which in any case is not required in our method, as we only compute rewards based on the final answer. One consequence is that if there is no model-generated correct solution for a question, then that question is not included in our training. We follow the same setup as before to first generate reasoning and then a final answer by the models (see subsection A.1 for prompt) to construct data for iterations of Iterative RPO. We only train on the training set (both easy and challenge sets) and do not utilize the ARC Corpus.
@@
We hence perform experiments using a very similar setup to the one previously described for GSM8K. Overall results are given in Table 2. We again find that Iterative RPO provides increased performance across iterations (84.8%, 86.2%, 86.7%) over three iterations. Majority voting using the model in the third iteration (32 samples, temperature 0.8) leads to another small boost (87.9%). These results outperform zero-shot CoT (77.8%), SFT on chosen sequences (79.8%) and standard DPO (83.5%).
@@ 3.3 MATH Task
We also experiment with more advanced math problems using the MATH (Hendrycks et al., 2021) dataset that is composed of 12,500 competition problems. The test set has 5,000 examples. Similar to the GSM8K dataset, a gold CoT solution is provided for each problem, and the gold answers can be matched uniquely to predicted answers after normalization to compute rewards. We do not use the accompanying pretraining data. For each MATH question, we use a few-shot prompt given in subsection A.1 as the input to the language model. In particular, the prompt includes four fixed in-context examples chosen from the training set. The language model needs these demonstrations so that the final answers can be properly formatted in .
@@
Results are given in Table 2. We again find that Iterative RPO provides increased performance across iterations, from 17.7% to 19.9% to 20.8% over three iterations. Majority voting (32 samples, temperature 0.8) leads to a significant boost in performance (29.1%). These results outperform few-shot CoT (12.5%), SFT on chosen sequences (16.8%) and standard DPO (12.4%). In particular, DPO degrades the performance compared to initialization.
Overall, we find on all three distinct tasks we tried, from simpler to more difficult, similar observations about the performance gains exhibited by our method.
@@ 4 Related Work
General Iterative Alignment Methods
@@
Methods Improving Reasoning Ability
While a number of approaches have been developed to curate or distill training data for reasoning tasks (Yu et al., 2023; Toshniwal et al., 2024), in this work we focus on learning algorithms which is an orthogonal axis. Expert Iteration assumes a reward model, and repeatedly uses rejection sampling to filter generations and train on them, which is found to match the sample complexity of PPO (Havrilla et al., 2024). STaR (Zelikman et al., 2022) relies on a similar loop: generate rationales to answer many questions, prompted with a few rationale examples; if the generated answers are wrong, try again to generate a rationale given the correct answer; and then fine-tune on all the rationales that ultimately yielded correct answers; and repeat. ReSTEM (Singh et al., 2024) assumes a ground truth verifier and also fine-tunes on filtered samples in a repeated fashion. All these methods rely on finding high-quality samples for SFT-like training, rather than using DPO-like pairwise preference optimization as in our work.
@@ Acknowledgments
Appendix A More Details on Experimental Setup
A.1 Prompts

### v3

@@ 3.2 ARC-Challenge Task
To test reasoning capabilities outside of mathematics, we employ ARC (Clark et al., 2018) which covers multiple science subjects. Questions are multiple-choice, for example: “A fold observed in layers of sedimentary rock most likely resulted from” with four possible answers, e.g., “(A) cooling of flowing magma, (B) converging of crustal plates, (C) deposition of river sediments, or (D) solution of carbonate minerals”. The training dataset contains 7.7k questions split into easy and challenge sets. We report results on the ARC-Challenge test set which has 1172 examples. There is no gold chain-of-thought reasoning provided for training examples in this task. Our method does not have that requirement and hence can still be applied as we only compute rewards based on the final answer. One consequence however is that if there is no model-generated correct solution for a question, then that question is not included in our training. We follow the same setup as before to first generate reasoning and then a final answer by the models (see subsection B.1 for prompt) to construct data for iterations of Iterative RPO. We only train on the training set (both easy and challenge sets) and do not utilize the supporting ARC Corpus.
@@
We hence perform experiments using a very similar setup to the one previously described for GSM8K. Overall results are given in Table 2. We again find that Iterative RPO provides increased performance across iterations (84.8%, 86.2%, 86.7%) over three iterations. Majority voting using the model in the third iteration (32 samples, temperature 0.8) leads to another small boost (87.9%). These results outperform zero-shot CoT (77.8%), SFT on chosen sequences (79.8%) and standard DPO (83.5%). We arrive at similar observations in 4(a) compared to Figure 3: when training with DPO without NLL loss, the log probabilities of chosen sequences barely increase over training; when training with DPO with NLL loss, the log probabilities increase noticeably.
@@
(a) ARC-Challenge
(b) MATH
Figure 4: Effect of NLL loss term on DPO training for ARC and MATH. The legend on the right plot is omitted due to space constraint, but it is the same as the legend in the left plot. Similar to GSM8K, in ARC-Challenge and MATH, we see that the log probabilities of chosen sequences barely increase over training steps when training with DPO. However, when training with DPO with NLL loss, the log probabilities increase over training steps.
@@ 3.3 MATH Task
We also experiment with more advanced math problems using the MATH (Hendrycks et al., 2021) dataset that is composed of 12,500 competition problems, for example the question: “Tom has a red marble, a green marble, a blue marble, and three identical yellow marbles. How many different groups of two marbles can Tom choose?”. While this may look superficially similar to the GSM8K task, it features substantially harder questions, as will be shown by the baseline performance. The test set has 5,000 examples. Similar to the GSM8K dataset, a gold CoT solution is provided for each problem, and the gold answers can be matched uniquely to predicted answers after normalization to compute rewards. We do not use the accompanying pretraining data. For each MATH question, we use a few-shot prompt given in subsection B.1 as the input to the language model. In particular, the prompt includes four fixed in-context examples chosen from the training set. The language model needs these demonstrations so that the final answers can be properly formatted in .
@@
Results are given in Table 2. We again find that Iterative RPO provides increased performance across iterations, from 17.7% to 19.9% to 20.8% over three iterations. Majority voting (32 samples, temperature 0.8) leads to a significant boost in performance (29.1%). These results outperform few-shot CoT (12.5%), SFT on chosen sequences (16.8%) and standard DPO (12.4%). In particular, DPO degrades the performance compared to initialization. Similar to the previous tasks, we show the log-probabilities during training in 4(b).
Overall, we find on all three distinct tasks we tried, from simpler to more difficult, similar observations on performance gains are exhibited by our method.
@@ 4 Related Work
General iterative alignment methods.
@@
Methods improving reasoning ability.
While a number of approaches have been developed to curate or distill training data for reasoning tasks (Yu et al., 2024; Toshniwal et al., 2024), in this work we focus on learning algorithms which is an orthogonal axis. Expert Iteration assumes a reward model, and repeatedly uses rejection sampling to filter generations and train on them, which is found to match the sample complexity of PPO (Havrilla et al., 2024). STaR (Zelikman et al., 2022) relies on a similar loop: generate rationales to answer many questions, prompted with a few rationale examples; if the generated answers are wrong, try again to generate a rationale given the correct answer; and then fine-tune on all the rationales that ultimately yielded correct answers; and repeat. ReSTEM (Singh et al., 2024) assumes a ground truth verifier and also fine-tunes on filtered samples in a repeated fashion. All these methods rely on finding high-quality samples for SFT-like training, rather than using DPO-like pairwise preference optimization as in our work.
@@ Acknowledgments
Appendix A Limitations
On experiments: When training iteration $t$ using iterative RPO, we do not make use of the collected data in previous iterations. Utilizing those data could potentially boost the performance even more. We leave this point to future work, as it is not central to our theses. In addition, we have experimented on three tasks. It is unclear how the approach would perform on general instruction tuning tasks without a clear best answer, but we argue that positive results on the three tasks in this paper can already prove the method useful. The current recipe requires correct answers, and a clear metric for comparing a generated response with this correct answer.
Regarding our loss function: The NLL loss is shown to be helpful in our case. Our iterative RPO algorithm requires training data to be mostly collected from the previous iteration of the model. Therefore, the chosen and rejected sequences all have reasonably high probability under the model distribution. When training sequences are arbitrary (e.g., sampled from other models), it is unclear whether the NLL loss is necessary (although this setting does not fall under the umbrella of the iterative RPO procedure in this paper).
Appendix B More Details on Experimental Setup
B.1 Prompts